on # 3. Ceja has utility function U=A2*B2 , where A equals the number of apples she eats each week, while B is the number of bananas she eats each week. Ceja has $20 to spend on fruit each week. The price of an apple is $1, while the price of a banana is $0.25. 

Find out the combination of Apples and Bananas that maximize Ceja’ satisfaction


1
Expert's answer
2021-05-24T13:41:55-0400

Maximizing utility condition is achieved where

The slope of indifference curve = slope of the budget constraint

"MRS =\frac{ P_

A}

{P_

B}"


"where MRS =\frac{ MU_

A

}{MU

_B}"


"MU\n\n_A\n\n = 2AB\n\n^2"

"MU\n\n_A\n\n = 2A^\n\n2\n\nB"

"Now MRS = \frac{P
_
A

}{P

_B}"


"\frac{MU_

A

}{MU

_B}



 = \frac{1

}{0.25}"


"\frac{2AB

^2

}{2A

^2

B}



 = \frac{1

}{0.25}"


"\frac{B}{

A}



 = 4"

"B = 4A"

Now Substitute the value of B= 4A in Budget constraint

"P\n\n_A\n\nA + P\n\n_B\n\nB = M"

"1A +0.25B = 20"

"1A +0.25\times4A = 20"

"1A +1A = 20"

"2A = 20"

"A = 10"

Now put A = 10 in budget constraint to calculate bundle B

"P\n\n_A\n\nA + P\n\n_B\n\nB = M"

"1\times 10+0.25B = 20"

"10+0.25B = 20"

"0.25B = 20 \u2212 10"

"0.25B = 10"

"B = \frac{10}{0.25}"

"B = 40"


therefore the combination of apple and banana to maximize the utility are (10,40); 10 apples and 40 bananas